OpenStudy (anonymous):

y/y^2+6y-7 * y^2+7y/y^3

OpenStudy (anonymous):

\[\frac{ y }{ y ^{2}+6y-7}\times \frac{ y ^{2}+7y }{ y ^{3} }\]